OpenStudy (janu16):

Hey can anyone help me? So I calculated the answer on a calculator but I wanna learn how to do this without calculator 6.5*10^4 + 6.2*10^5

OpenStudy (greatlife44):

so it's actually easy. whenever you multiply a number by 10 you add a zero. 6*10 = 60, add one zero. 6*100 =600 add two zeros

OpenStudy (greatlife44):

\[10^{4} = 10,000\] 10,000*6.5 = ?

OpenStudy (janu16):

685000

OpenStudy (janu16):

I mean 65000

OpenStudy (greatlife44):

good

OpenStudy (greatlife44):

\[10^{5} = 100,000\] \[100,000*6.2 = ? \]

OpenStudy (janu16):

620000 than you would add those and get 685000 right?

OpenStudy (greatlife44):

so then you would add the two together: \[620,000+65,000 = 685,000\]

OpenStudy (janu16):

Thx

OpenStudy (triciaal):

the objective is to learn exponents and scientific notation design to do calculations with very large and very small numbers.

OpenStudy (triciaal):